# An Empirical Risk Minimization Approach for Offline Inverse RL and Dynamic Discrete Choice Model

### v4

@@ 3. Deriving the PL Condition
Recall that we say $f$ satisfies $\mu$-$PL$ condition if $2\mu(f(\bm{\theta})-f\left(\bm{\theta}^{\ast}\right))\leq\|\nabla f(\bm{\theta})\|^{2}$.
@@
Now suppose, for the purpose of contradiction, that there exists $\hat{\bm{\theta}}\in\Theta$ such that $\left\langle\nabla f_{1}(\hat{\bm{\theta}}),\nabla f_{2}(\hat{\bm{\theta}})\right\rangle<0$. Consider the direction $d:=-g_{1}=-\nabla f_{1}(\hat{\bm{\theta}})$. Then $\nabla f_{1}(\hat{\bm{\theta}})^{\top}d=g_{1}^{\top}\left(-g_{1}\right)=-\left\|g_{1}\right\|^{2}<0$ holds. This implies that $f_{1}(\hat{\bm{\theta}}+\eta d)<f_{1}(\hat{\bm{\theta}})$. Then QG condition for $f_{1}$ implies that
@@
Now, note that $\nabla f_{2}(\hat{\bm{\theta}})^{\top}d=g_{2}^{\top}\left(-g_{1}\right)=-g_{1}^{\top}g_{2}$. Since $g_{1}^{\top}g_{2}<0$, $\nabla f_{2}(\hat{\bm{\theta}})^{\top}d>0$. Therefore, $f_{2}(\hat{\bm{\theta}}+\eta d)>f_{2}(\hat{\bm{\theta}})$ for sufficiently small $\eta>0$. That is, $f_{2}(\hat{\bm{\theta}}+\eta d)-f_{2}(\bm{\theta}^{\ast})>f_{2}(\hat{\bm{\theta}})-f_{2}(\bm{\theta}^{\ast})$. By QG condition, this implies that $\left\|\hat{\bm{\theta}}+\eta d-\bm{\theta}^{*}\right\|>\left\|\hat{\bm{\theta}}-\bm{\theta}^{*}\right\|$. Contradiction.
@@ 1. Optimization error analysis
Applying Algorithm 1 using stochastic gradient descent (SGD), $P_{T}$ satisfies
@@
$\mu_{1}\cdot\left(\hat{\bm{\theta}}_{T}-\hat{\bm{\theta}}_{\mathcal{D}_{N}}\right)^{2}\leq\frac{\nu}{\gamma+T}.$
@@
Therefore, we can conclude that $\cdot\left(\hat{\bm{\theta}}_{T}-\hat{\bm{\theta}}_{\mathcal{D}_{N}}\right)^{2}\leq\mathcal{O}(1/T)$, or $|\hat{\bm{\theta}}_{T}-\hat{\bm{\theta}}_{\mathcal{D}_{N}}|$ is $\mathcal{O}(1/\sqrt{T})$.
2. Statistical error analysis.
Throughout, we closely follow Charles and Papailiopoulos (2018). Note that $\beta\in(0,1)$, $\|r\|_{\infty}<R_{\max}$ for some $R_{\max}<\infty$, and the Bellman equation holds. Therefore, there exist $Q_{\min},Q_{\max}\in\mathbb{R}$ such that
Consequently, $V_{Q_{\theta}}\left(s^{\prime}\right)=\log\sum_{b}\exp\left(Q_{\theta}\left(s^{\prime},b\right)\right)$ satisfies $\log\left(|\mathcal{A}|e^{Q_{\text{min }}}\right)\leq V_{Q_{\theta}}\left(s^{\prime}\right)\leq\log\left(|\mathcal{A}|e^{Q_{\text{max }}}\right)$. Let $\left\|V_{Q_{\theta}}\right\|_{\infty}\leq V_{\max}<\infty$. Define $M=:\left(R_{\max}+\beta V_{\max}+Q_{\max}\right)$. Also, from Assumption 5, let $\left\|\nabla_{\theta}Q_{\theta}\right\|_{2}\leq B$, Then by Lemma 28, $\mathcal{R}_{\text{emp}}$ is Lipschitz continuous in $\bm{\theta}$ with constant $L:=(1+2M(1+\beta))B$.
Denote the minimizer of expected risk function $\mathcal{R}_{\text{exp}}$ as $\bm{\theta}^{\ast}$ and the minimizer of empirical risk function $\mathcal{R}_{\text{emp}}$ for the data set $\mathcal{D}_{N}$ as ${\bm{\theta}}^{\ast}_{{\mathcal{D}}_{N}}$, where $|\mathcal{D}_{N}|=N$. Then by Theorem 3.5 of Charles and Papailiopoulos (2018) and Lemma 5 of Lei (2023), $\mathcal{R}_{\text{emp}}$ being PL in $\bm{\theta}$ implies that ${\bm{\theta}}^{\ast}_{\mathcal{D}_{N}}$ and $\hat{\bm{\theta}}^{\ast}$ satisfies
@@
$\bigl{|}\mathbb{E}_{\mathcal{D}_{N}}\bigl{[}\mathcal{R}_{exp}(\bm{\theta}^{\ast}_{\mathcal{D}_{N}})-\mathcal{R}_{emp}(\bm{\theta}^{\ast}_{\mathcal{D}_{N}};\mathcal{D}_{N})\bigr{]}\bigr{|}\leq\frac{2L^{2}}{\mu_{1}N}.$
@@
$\bigl{|}\mathbb{E}_{\mathcal{D}_{N}}\bigl{[}\mathcal{R}_{exp}(\bm{\theta}^{\ast})-\mathcal{R}_{emp}(\bm{\theta}^{\ast};\mathcal{D}_{N})\bigr{]}\bigr{|}\leq\frac{2L^{2}}{\mu_{1}N}.$
@@
$\mathcal{R}_{\exp}\left(\bm{\theta}_{D_{N}}^{*}\right)-\mathcal{R}_{\exp}\left(\bm{\theta}^{*}\right)$
$=\left[\mathcal{R}_{\exp}\left(\bm{\theta}_{D_{N}}^{*}\right)-\mathcal{R}_{\text{emp }}\left(\bm{\theta}_{D_{N}}^{*};\mathcal{D}_{N}\right)\right]+\underbrace{\left[\mathcal{R}_{\text{emp }}\left(\bm{\theta}_{\mathcal{D}_{N}}^{*};\mathcal{D}_{N}\right)-\mathcal{R}_{\text{emp}}\left(\bm{\theta}^{*};\mathcal{D}_{N}\right)\right]}_{\leq 0}+\left[\mathcal{R}_{\text{emp }}\left(\bm{\theta}^{*};\mathcal{D}_{N}\right)-\mathcal{R}_{\text{exp }}\left(\bm{\theta}^{*}\right)\right]$
We have
$\mathbb{E}_{\mathcal{D}_{N}}\bigl{[}\mathcal{R}_{\exp}\left(\bm{\theta}_{\mathcal{D}_{N}}^{*}\right)-\mathcal{R}_{\exp}\left(\bm{\theta}^{*}\right)\bigr{]}\leq\frac{4L^{2}}{\mu_{1}N}.$
Since being of $\mathcal{C}^{2}$ and PL implies Quadratic growth (QG) condition (Rebjock and Boumal 2023), we have
$\mathbb{E}_{\mathcal{\mathcal{D}_{N}}}\left[\left(\bm{\theta}^{\ast}_{\mathcal{D}_{N}}-\bm{\theta}^{\ast}\right)^{2}\right]\leq\frac{1}{\mu_{1}}\cdot\frac{4L^{2}}{\mu_{1}N}=\frac{4((1+2M(1+\beta))B)^{2}}{\mu_{1}^{2}N}$

### v6

@@ 3. Deriving the PL Condition
Recall that we say $f$ satisfies $\mu$-PLPL condition if $2\mu(f(\bm{\theta})-f\left(\bm{\theta}^{\ast}\right))\leq\|\nabla f(\bm{\theta})\|^{2}$.
@@
[…]
@@
[…]
@@ 1. Optimization error analysis
Applying Algorithm 1 using stochastic gradient ascent descent (SGAD), $P_{T}$ satisfies
@@
$\mu_{1}\cdot\text{dist}^{2}\left(\hat{\bm{\theta}}_{T},S_{\mathcal{D}_{N}}\right)\leq\frac{\nu}{\gamma+T}.$
@@
where $S_{\mathcal{D}_{N}}$ is the set of minimizers of empirical risk, and $\text{dist}\left({\bm{\theta}},S\right):=\min_{\bm{\theta}^{\prime}\in S}\|\bm{\theta}-\bm{\theta}^{\prime}\|$. Therefore, we can conclude that $\text{dist}^{2}\left(\hat{\bm{\theta}}_{T},S_{\mathcal{D}_{N}}\right)\leq\mathcal{O}(1/T)$, or $\text{dist}\left(\hat{\bm{\theta}}_{T},S_{\mathcal{D}_{N}}\right)$ is $\mathcal{O}(1/\sqrt{T})$.
2. Global convergence analysis.
The challenge of statistical convergence analysis using data from a Markov Chain, which is non-i.i.d sampled data, is well known (Mohri and Rostamizadeh 2010; Fu et al. 2023; Abeles et al. 2024). The following Lemma 28 by Kang and Jang (2025) shows that applying Algorithm 1 for Bellman residual minimization has sample complexity of $\mathcal{O}(1/N)$.
Lemma 28 (Sample complexity of Bellman residual minimization using SGDA (Kang and Jang 2025)).
$\bigl(\widehat{\bm{\theta}}_{1}^{(T)},\widehat{\bm{\theta}}_{2}^{(T)}\bigr)$ be the parameters returned by Algorithm 1 after $T$ SGDA iterations on the empirical objective (15) with stepsizes $\eta_{t}=\frac{c_{1}}{c_{2}+T}$. Define the population risk
$\mathcal{R}(\bm{\theta}_{2})$ $:=\mathbb{E}_{(s,a)\sim\pi^{\ast},\nu_{0}}\,\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\!\Bigl[\mathcal{L}_{\mathrm{TD}}\!\bigl(Q_{\bm{\theta}_{2}}\bigr)(s,a,s^{\prime})-\beta^{2}\!\bigl(V_{Q_{\bm{\theta}_{2}}}(s^{\prime})-\zeta_{\widetilde{\bm{\theta}}_{1}^{\star}}(s,a)\bigr)^{2}\Bigr],$
$\widehat{\mathcal{R}}_{n}(\bm{\theta}_{2})$ $:=\frac{1}{N}\sum_{(s,a,s^{\prime})\in\mathcal{D}_{\pi^{\ast},\nu_{0}}}\Bigl[\mathcal{L}_{\mathrm{TD}}\!\bigl(Q_{\bm{\theta}_{2}}\bigr)(s,a,s^{\prime})-\beta^{2}\!\bigl(V_{Q_{\bm{\theta}_{2}}}(s^{\prime})-\zeta_{\widetilde{\bm{\theta}}_{1}^{\star}}(s,a)\bigr)^{2}\Bigr],$
where $\widetilde{\bm{\theta}}_{1}^{\star}$ is the minimizer in
$\widetilde{\bm{\theta}}_{1}^{\star}\in\operatorname{argmin}_{\bm{\theta}_{1}\in\bm{\Theta}}\frac{1}{N}\sum_{\left(s,a,s^{\prime}\right)\in\mathcal{D}_{\pi^{\ast},\nu_{0}}}\left[\left(V_{Q_{\bm{\theta}_{2}}}\left(s^{\prime}\right)-\zeta_{\bm{\theta}_{1}}(s,a)\right)^{2}\right]$
Let $\bm{\Theta}_{2}^{\ast}:=\operatorname{argmin}_{\bm{\theta}_{2}\in\bm{\Theta}}\mathcal{R}(\bm{\theta}_{2})$ denote the (possibly non-singleton) set of population minimizers, and fix any $\bm{\theta}_{2}^{\ast}\in\bm{\Theta}_{2}^{\ast}$. Then
$\mathbb{E}\!\Bigl[\mathcal{R}\!\bigl(\widehat{\bm{\theta}}_{2}^{(T)}\bigr)-\widehat{\mathcal{R}}_{n}\!\bigl(\widehat{\bm{\theta}}_{2}^{(T)}\bigr)\Bigr]\;\leq\;(1+L/\rho)\,G\,\varepsilon_{T},$
@@
$\mathbb{E}\Bigl[\mathcal{R}\!\bigl(\widehat{\bm{\theta}}_{2}^{(T)}\bigr)-\mathcal{R}\!\bigl(\bm{\theta}_{2}^{\ast}\bigr)\Bigr]\;\leq\;(1+L/\rho)G\varepsilon_{T}+\frac{\nu}{\gamma+T}$
where
$\varepsilon_{T}\;=\;O\!\left((c_{2}+T)^{-\min\left\{\tfrac{1}{2},\;\tfrac{3c\,c_{1}}{8}\right\}}\right)\;+\;\!\frac{C}{N}$
Here, $L$ is the Lipschitz smoothness constant from Lemma 29, $\rho$ is the strongly concavity constant, $G$ is the upper bound on the norm of the stochastic gradients, $c_{1}$ and $c_{2}$ are from step-size choice, $c$ is $\min(\mu_{PL}/2,\rho/2)$, and $C$ is a constant defined in Kang and Jang (2025).
Using $\mathbb{E}\Bigl[\mathcal{R}\!\bigl(\widehat{\bm{\theta}}_{2}^{(T)}\bigr)-\mathcal{R}\!\bigl(\bm{\theta}_{2}^{\ast}\bigr)\Bigr]\;\leq\;(1+L/\rho)G\varepsilon_{T}+\frac{\nu}{\gamma+T}$ and the fact that $\mathcal{R}$ satisfies PL condition, which is equivalent to QG condition under $\mathcal{C}^{2}$, we arrive at the final global convergence analysis result (to the solution set):
@@
$\mathbb{E}\Bigl[\operatorname{dist}\!\bigl(\widehat{\bm{\theta}}_{2}^{(T)},\bm{\Theta}_{2}^{\ast}\bigr)^{2}\Bigr]$ $\leq\frac{1}{\mu_{PL}}\mathbb{E}\Bigl[\mathcal{R}\!\bigl(\widehat{\bm{\theta}}_{2}^{(T)}\bigr)-\mathcal{R}\!\bigl(\bm{\theta}_{2}^{\ast}\bigr)\Bigr]\;\leq\;(\frac{1}{\mu_{PL}}+\frac{L}{\rho\mu_{PL}})G\varepsilon_{T}+\frac{1}{\mu_{PL}}\frac{\nu}{\gamma+T}$
$=O\!\left((c_{2}+T)^{-\min\left\{\tfrac{1}{2},\;\tfrac{3c\,c_{1}}{8}\right\}}\right)\;+\;\!\frac{C^{\prime}}{N}$
for some $C^{\prime}$, where $\operatorname{dist}(\theta,S):=\min_{\theta^{\prime}\in S}\|\theta-\theta^{\prime}\|_{2}$.
Next, we convert the above bound into the bound in terms of $\mathbb{E}_{\mathcal{D}_{N}}\left[\left\|Q_{\bm{\theta}_{2}^{\star}}-Q_{\widehat{\bm{\theta}}_{2}^{(T)}}\right\|_{\infty}^{2}\right]$. The key idea is to utilize $\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left\|\nabla_{\theta}Q_{\theta}(s,a)\right\|_{2}\leq B$ from Assumption 5.
Fix any $\theta_{1},\theta_{2}\in\Theta$ and any $(s,a)$. Consider the line segment
$\theta(t)=\theta_{2}+t\left(\theta_{1}-\theta_{2}\right),\quad t\in[0,1].$
By the fundamental theorem of calculus (equivalently, the mean value theorem in integral form),
$Q_{\theta_{1}}(s,a)-Q_{\theta_{2}}(s,a)=\int_{0}^{1}\nabla_{\theta}Q_{\theta(t)}(s,a)^{\top}\left(\theta_{1}-\theta_{2}\right)dt$
Take absolute values and apply Cauchy-Schwarz:
$\left|Q_{\theta_{1}}(s,a)-Q_{\theta_{2}}(s,a)\right|\leq\int_{0}^{1}\left\|\nabla_{\theta}Q_{\theta(t)}(s,a)\right\|_{2}\left\|\theta_{1}-\theta_{2}\right\|_{2}dt\leq\int_{0}^{1}B\left\|\theta_{1}-\theta_{2}\right\|_{2}dt=B\left\|\theta_{1}-\theta_{2}\right\|_{2}$
where the both inequalities are from $\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left\|\nabla_{\theta}Q_{\theta}(s,a)\right\|_{2}\leq B$ of Assumption 5. Now take supremum over $(s,a)$ :
$\left\|Q_{\theta_{1}}-Q_{\theta_{2}}\right\|_{\infty}:=\sup_{(s,a)}\left|Q_{\theta_{1}}(s,a)-Q_{\theta_{2}}(s,a)\right|\leq B\left\|\theta_{1}-\theta_{2}\right\|_{2}$
So the map $\theta\mapsto Q_{\theta}$ is $B$-Lipschitz from $\left(\Theta,\|\cdot\|_{2}\right)$ into $\left(\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R},\|\cdot\|_{\infty}\right)$. Therefore, for any $\theta^{*}\in\Theta_{2}^{*}$,
$\left\|Q_{\widehat{\bm{\theta}}_{2}^{(T)}}-Q_{\theta^{*}}\right\|_{\infty}\leq B\left\|\widehat{\bm{\theta}}_{2}^{(T)}-\theta^{*}\right\|_{2}$
Take $\inf$ over $\theta^{*}\in\Theta_{2}^{*}$ on both sides:
$\inf_{\theta^{*}\in\Theta_{2}^{*}}\left\|Q_{\widehat{\bm{\theta}}_{2}^{(T)}}-Q_{\theta^{*}}\right\|_{\infty}\leq B\inf_{\theta^{*}\in\Theta_{2}^{*}}\left\|\widehat{\bm{\theta}}_{2}^{(T)}-\theta^{*}\right\|_{2}=B\operatorname{dist}\left(\widehat{\bm{\theta}}_{2}^{(T)},\Theta_{2}^{*}\right)$
@@
$Q_{\theta^{*}}\equiv Q^{*}\quad\text{ for all }\theta^{*}\in\Theta_{2}^{*},$
then the set $\left\{Q_{\theta^{*}}:\theta^{*}\in\Theta_{2}^{*}\right\}$ is a singleton $\left\{Q^{*}\right\}$. Therefore,
$\inf_{\theta^{*}\in\Theta_{2}^{*}}\left\|Q_{\widehat{\bm{\theta}}_{2}^{(T)}}-Q_{\theta^{*}}\right\|_{\infty}=\left\|Q_{\widehat{\bm{\theta}}_{2}^{(T)}}-Q^{*}\right\|_{\infty}.$
Therefore, we have
$\left\|Q^{\ast}-Q_{\widehat{\bm{\theta}}_{2}^{(T)}}\right\|_{\infty}\;\leq\;B\,\operatorname{dist}\!\bigl(\widehat{\bm{\theta}}_{2}^{(T)},\bm{\Theta}_{2}^{\ast}\bigr).$
Take expectations and squares:
$\mathbb{E}_{\mathcal{D}_{N}}\left[\left\|Q^{\ast}-Q_{\widehat{\bm{\theta}}_{2}^{(T)}}\right\|_{\infty}^{2}\right]\leq B^{2}\mathbb{E}_{\mathcal{D}_{N}}\left[\operatorname{dist}\!\bigl(\widehat{\bm{\theta}}_{2}^{(T)},\bm{\Theta}_{2}^{\ast}\bigr)^{2}\right].$
Therefore, we achieve
$\mathbb{E}_{\mathcal{D}_{N}}\left[\left\|Q^{\ast}-Q_{\widehat{\bm{\theta}}_{2}^{(T)}}\right\|_{\infty}^{2}\right]\leq B^{2}\left[O\left(\left(c_{2}+T\right)^{-\min\left\{\frac{1}{2},\frac{3cc_{1}}{8}\right\}}\right)+\frac{C^{\prime}}{N}\right].$